OpenStudy (anonymous):

Show that 2/5 √2 is an irrational number (I)

OpenStudy (unklerhaukus):

your question isn't clear

OpenStudy (unklerhaukus):

\[\frac 25 \sqrt 2\]

OpenStudy (anonymous):

Verifies that \[\frac{ 2 }{ 5 }\sqrt{2}\] is the irrational number.